It takes James 55 seconds to run around a 400 meter circular track. It takes Henry 62 seconds to run around the same track. If they started at the same point on the track but ran in opposite directions how long would it take for them to meet each other.

Answer by ankor@dixie-net.com(22740)

It takes James 55 seconds to run around a 400 meter circular track.
It takes Henry 62 seconds to run around the same track.
If they started at the same point on the track but ran in opposite directions how long would it take for them to meet each other.
:
let t = time for them to meet
:
Find their speed in meter/sec
James: 400/55 = 7.27 m/sec
Henry: 400/62 = 6.45 m/sec
:
When they meet the total distance they have run = 400 m
:
Write a distance equation, dist = speed * time
7.27t + 6.45t = 400
13.72t = 400
t = 400/13.72
t = 29.15 sec for them to meet